Use either the fundamental counting principle or the permutation formulas (or both) to answer the following
How many seven-digit phone numbers can be formed if the first digit cannot be 0 and
repetition of digits is not permitted?
if repetition of digits is permitted?
The Board of Directors does not have assigned


Answer by stanbon(75887)

Use either the fundamental counting principle or the permutation formulas (or both) to answer the following
How many seven-digit phone numbers can be formed if the first digit cannot be 0
and
repetition of digits is not permitted?
Ans:
1st digit: 9 ways
2nd digit: 9 ways
3rd to 7th digits: 8!/2! = 8*7*6*5*4*3 = 20160
Total 9^2*20160 = 1632960
--------------------------
if repetition of digits is permitted?
Ans: 9*10^6 = 9 million
--------------------------
Cheers,
Stan H.